OpenStudy (anonymous):

Really need help with some algebra 2

OpenStudy (anonymous):

Thank you. Hold on one second please

OpenStudy (anonymous):

Given that f(x)=-x+4 and g(x)=-2x-3 solve for f(g(x) when x=2

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

and after this one i have a few more questions if you dont mind

OpenStudy (anonymous):

im sorry no. Im confused lol

OpenStudy (anonymous):

i got -5 as the answer but that isnt an option

OpenStudy (anonymous):

are you there?